Please Help. Im having trouble with these equations.
1.

A. √x-2=1


B. √xˆ3=27

C. 3√xˆ2=9

Answer by checkley71(8403)

IF THIS IS THE PQOBLEM SQRT(X-2)=1 THEN SQRT1=1 THUS X-2=1 OR X=1+2 OR X=3
PROOF
SQRT(3-2)=1
SQRT1=1
1=1
----------------------------------------
SQRTX^3=27 OR X^3=27*27 X^3=3*3*3*3*3*3 OR X^3=9*9*9 X=9
SQRT9^3=27
SQRT9*9*3*3=27
9*3=27
27=27
--------------------------------
3SQRTX^2=9
3X^2=9*9 OR X^2=81/3 X^2=27 X^2=3*3*3 X=3SQRT3
----------------------------------------------------